OpenStudy (anonymous):

T or F?. log[1]512 is an expression that CANNOT be evaluated?

OpenStudy (jamesj):

Building on what you've been taught already, if this expression could be evaluated, then there would be a number n such that 1^n = 512 Is there such a number, n?

OpenStudy (anonymous):

would this also apply for n like the last question?

OpenStudy (jamesj):

It's a different question, so evaluate this one on its face. Let's experiment for a moment: 1^2 = 1 1^200 = 1 1^100000000000000000000 = 1 so making n big doesn't work 1^(-1) = 1 1^(-100000000000000) = 1 nor does making it large and negative or small ... hmm ... 1^(1/2) = 1 1^(1/100000000000) = 1 and 1^(-1/10000000000) = 1

OpenStudy (jamesj):

So do we think there is an n such that \[1^n = 512\] ?

OpenStudy (jamesj):

If there were such a number n, then log[1](512) = n.

OpenStudy (jamesj):

But .......